Write an equation of the line that is the perpendicular bisector of the line segment having endpoints (3,-1) and (3,5).
Answer by stanbon(75887)

Write an equation of the line that is the perpendicular bisector of the line segment having endpoints (3,-1) and (3,5).
------
slope of line thru given points: (5--1)/(3-3) = undefined
----
Therefore slope of perpendicular bisector: m = 0
-------------------
Midpoint of given segment:
x = (3+3)/2 = 3
y = (-1+5)/2 = 2
----
Midpoint is (3,2)
---
If slope = 0, the line is a horizontal line thru (3,2)
---
Equation of perp/bisector: y = 2
======================================
Cheers,
Stan H.
============================